OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

y=mx+b. First find slope of both those points and that = m. plug in one of those points to find the value for b. once you have a complete y=mx+b equation you can then move x and y values to one side and make it standard form.

OpenStudy (anonymous):

Let (2, 2) and (6, 3) be (x1,y1) & (x2,y2) First find the slope m=(y2-y1)/(x2-x1) m=1/4 Then take any point (2, 2) or (6, 3) (y-y1)=m(x-x1) Then substitute the values of m & x1&y1 x-4y+6=0

OpenStudy (anonymous):

so which one is that the last option ?

OpenStudy (anonymous):

Yes.